OpenStudy (ray10):

can't identify what I'm missing: anti-differentiate 4x*(SQRT)x^2 +3 dx

sam (.sam.):

You mean \[\int\limits 4x \sqrt{x^2+3}dx\]

OpenStudy (ray10):

\[\int\limits_{}^{} 4x \sqrt{x ^{2}+3} dx\]

OpenStudy (ray10):

yup! I've been trying to work it out and it comes to the end but

sam (.sam.):

Have you tried using u-sub \[u=x^2+3\]

OpenStudy (ray10):

yes I subbed that in as u, and got \[\frac{ du }{ dx } = 4x\] the answer is \[\frac{ 4 }{ 3 }(x ^{2}+3)^{\frac{ 3 }{ 2 }}+c\]

OpenStudy (ray10):

when I work it out after coming to the simplified equation of \[\int\limits_{}^{} \sqrt{u} du\]

sam (.sam.):

Do it like this, \[u=x^2+3\] \[du=2xdx\] \[\frac{du}{2}=xdx\] ------------------------------------------ \[\int\limits\limits 4x \sqrt{x^2+3}dx\] \[4\int\limits\limits \sqrt{u}\frac{du}{2}\] \[2 \int\limits \sqrt{u}du\]

sam (.sam.):

You forgot the 2

OpenStudy (ray10):

oh gosh, I just realized I made \[u = x ^{2} +3 = 4x \] >.<

OpenStudy (ray10):

thank you! so in questions like this remove the constant outside the integral?

OpenStudy (anonymous):

yes

OpenStudy (ray10):

ah alright thank you, do I nee to close this question to ask another question or can I just ask it here?

OpenStudy (ray10):

need*